OpenStudy (anonymous):

HELP!? Green’s Grocery sampled 730 shoppers to determine their interest in buying fresh fish. Of the people surveyed, 222 said they would buy fresh fish. Green’s Grocery estimates that 82,000 people visit the store at least once every two weeks and that people buying fresh fish would buy some at every visit. What can Green's Grocery estimate to be the annual fish sales rounded to the nearest ten thousand?

OpenStudy (tkhunny):

Kind of a pathetic question. Not your fault. Proportion: 222/730 = 0.30410958904109589041095890410959 So 30%-ish 0.30 * 82000 = 24600 Not every week, but every other week. 24600 * (52/2) = 639600 fish sales per year That's a pretty weak prediction. We haven't even discussed the size of the order.

OpenStudy (anonymous):

size of what order?

OpenStudy (tkhunny):

If you buy one fish stick, does that count? If you buy 300 pounds of red snapper. Is that an order? Those are amazingly different transaction! Anyway, we were not asked to worry about that. Did you follow the weak approximation?

OpenStudy (anonymous):

I'm lost? this class I'm in is so confusing. I follow the steps they give me and it doesn't end up right.

OpenStudy (ranga):

I am getting: 82000 x 222 / 730 x 52 /2 = 648361 and when rounded to the nearest ten thousand it is: 650,000

OpenStudy (anonymous):

I just checked that and it's right! thank you for showing me how!